OpenStudy (anonymous):

Select the pair of equations whose graphs are perpendicular. 2x - 8y = 9 12x - 3y = 7 3x + 6y = 8 y = 2x - 8 y = 2x - 7 x + y = 3 2y =- 3x + 5 2x + 3y = 4

OpenStudy (anonymous):

I hear it's really cool when you solve for \(y\).

OpenStudy (hoblos):

the product of slopes of perpendicular lines is -1 3x + 6y = 8 ----> m=-3/6=-1/2 y = 2x - 8 -----> m'=2 m*m' = -1 3x + 6y = 8 and y = 2x - 8 are perpendicular